OpenStudy (anonymous):

in a town election, candidates A and B were running for mayor. there were 30,500 people eligible to vote, and 80% of them actually voted. candidate A recieved 6100 votes. What percent of the votes cast did candidate B receive in a town election, candidates A and B were running for mayor. there were 30,500 people eligible to vote, and 80% of them actually voted. candidate A recieved 6100 votes. What percent of the votes cast did candidate B receive @Mathematics

OpenStudy (anonymous):

75 %

OpenStudy (anonymous):

show a work

OpenStudy (anonymous):

plz show how to solve the problem

OpenStudy (anonymous):

answer is 75%

OpenStudy (anonymous):

30500*80/100=24400 gave vote.... A got 6100....so B got 24400-6100=18300 votes....so the percentage is 18300*100/30500=60

OpenStudy (anonymous):

ohh...sorry i missed the last sentence ..18300*100/24400=75

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

welcome :)